# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ Acknowledgments
This work is supported generously by the NSF grants CCF-2312204 and CCF-2312205 and Air Force Office of Scientific Research FA9550-20-1-0397. Additional support is gratefully acknowledged from NSF 1915967, 2118199, and 2229012.
@@ Appendix A Formulations of Distributionally Robust Stochastic Control
[…]
@@ A.1 Admissible Policies
Notice that in this model, the history-dependent controller cannot directly use the realized action $w_{t}$ of the adversary to make its decision. This should be compared to the stochastic game settings González-Trejo et al. [2002] in which either player observes the action of the other and makes a decision based on such observation. However, this model can include the settings for which both players see the action of each other by considering a new state process $z_{t}=(x_{t},w_{t-1})$ and defining the state space and histories using $z_{t}$ instead of $x_{t}$.
@@ A.2 The Distributionally Robust Stochastic Control Problem
$v^{*}(\mu,\Pi,\Gamma):=\sup_{\pi\in\Pi}\inf_{\gamma\in\Gamma}E_{\mu}^{\pi,\gamma}\sum_{t=1}^{\infty}\alpha^{t}r(X_{t},A_{t})$ (A.3)
@@ A.3 Dynamic Programming
In this section, we show that the solutions to the distributional robust Bellman equations (2.2) and (2.3) will correspond to the DRSC value (A.3).
@@
A history-dependent version of Theorem 7 is established in González-Trejo et al. [2002]. In this paper, we will prove the more technically interesting Theorem 8. The proof of the rest of the Theorem 7 can be easily achieved by adapting the same proof techniques to deal with continuous state spaces in this paper to that of Theorem 2 from Wang et al. [2023c].
@@
B.1 Proof of Proposition 2.1
We prove Proposition 2.1 by applying the Banach fixed-point to the mapping $\mathcal{T}$ and $\overline{\mathcal{T}}$.
@@
Hence, $\left\|u^{\prime}\right\|\leq\beta\left\|u^{\prime}\right\|$.
@@ B.1.1 Proof of Lemma 2
$\leq\sup_{\phi\in\mathcal{Q}}\left|\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}r(x,a)+\alpha u(f(x,a,w))\phi\times\psi(da,dw)+\sup_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}-r(z,a)-\alpha u(f(z,a,w))\phi\times\psi(da,dw)\right|$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\sup_{\phi\in\mathcal{Q}}\sup_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}|r(x,a)-r(z,a)|+\alpha|u(f(x,a,w))-\alpha u(f(z,a,w))|\phi\times\psi(da,dw)$
$\leq\delta+\sup_{a\in\mathbb{A},w\in\mathbb{W}}\sup_{y\in\bar{B}_{f(x,a,w)}(\delta^{\prime})}\alpha|u(f(x,a,w))-u(y)|$
@@ B.2.1 Proof of Auxiliary Results for Theorem 8
Recall from 2.1 that $\left\|\bar{u}^{*}\right\|\leq\beta r_{\vee}$. Define and denote the $T$-step truncated value with terminal reward $\bar{u}^{*}$ by
$v_{T}(x,\pi,\gamma):=E_{x}^{\pi,\gamma}\left[\sum_{t=0}^{T_{\epsilon}-1}\alpha^{t}r(X_{t},A_{t})+\alpha^{T_{\epsilon}}\bar{u}^{*}(X_{T_{\epsilon}})\right].$ (B.2)
@@
$\left|v(x,\pi,\gamma)-v_{T_{\eta}}(x,\pi,\gamma)\right|$ $=\left|E_{x}^{\pi,\gamma}\left[\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},A_{t})\right]-E_{x}^{\pi,\gamma}\left[\sum_{t=0}^{T_{\eta}-1}\alpha^{t}r(X_{t},A_{t})+\alpha^{T_{\eta}}\bar{u}^{*}(X_{T_{\eta}})\right]\right|$ (B.3)
$=\left|E_{x}^{\pi,\gamma}\left[\sum_{t=T_{\eta}}^{\infty}\alpha^{t}r(X_{t},A_{t})-\alpha^{T_{\eta}}\bar{u}^{*}(X_{T_{\eta}})\right]\right|$
$\leq\alpha^{T_{\eta}}\left(\left|E_{x}^{\pi,\gamma}\sum_{t=T_{\eta}}^{\infty}\alpha^{t-T_{\eta}}r(X_{t},A_{t})\right|+\left|E_{x}^{\pi,\gamma}\bar{u}^{*}(X_{T_{\eta}})\right|\right)$
@@
$\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})$
@@
$\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta.$ (B.4)
@@
$\left\{{\psi\in\mathcal{P}:\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta}\right\}$
@@
$\left\{{h_{t}\in\mathbf{H}_{t}:\varnothing\neq\left\{{\psi\in\mathcal{P}:\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta}\right\}\cap\mathcal{D}}\right\}\in\mathcal{B}(\mathbf{H}_{t}).$
@@
$\iff$ $\exists\psi\in\mathcal{D}:\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta$
$\iff$ $\inf_{\psi\in\mathcal{D}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)-\bar{u}^{*}(x_{t})-\delta\leq 0$
@@
$\lim_{k\rightarrow\infty}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w_{k}))\pi_{t}(da|h_{t})=\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})$
i.e. $w\rightarrow\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w_{k}))\pi_{t}(da|h_{t})$ is bounded continuous. Hence,
$\psi\rightarrow\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)$
@@
$c(h_{t}):=\inf_{\psi\in\mathcal{D}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)-\bar{u}^{*}(x_{t})-\delta$
@@
$\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\bar{\gamma}^{\delta}_{t}(dw|h_{t})\leq\bar{u}^{*}(x_{t})+\delta.$
@@
$E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{t},A_{t})+\alpha\bar{u}^{*}(X_{t+1})]$ $=E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{t},A_{t})+\alpha\bar{u}^{*}(f(X_{t},A_{t},W_{t}))]$
$=E_{x}^{\pi,\bar{\gamma}^{\delta}}E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{t},A_{t})+\alpha\bar{u}^{*}(f(X_{t},A_{t},W_{t}))|H_{t}]$
$=E_{x}^{\pi,\bar{\gamma}^{\delta}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(X_{t},a)+\alpha\bar{u}^{*}(f(X_{t},a,w))\pi_{t}(da|H_{t})\bar{\gamma}^{\delta}_{t}(dw|H_{t})$
@@
$v_{T}(x,\pi,\bar{\gamma}^{\delta})$ $=E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\sum_{t=0}^{T-1}\alpha^{t}r(X_{t},A_{t})+\alpha^{T}\bar{u}^{*}(X_{T})\right]$ (B.5)
$=E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\sum_{t=0}^{T-2}\alpha^{t}r(X_{t},A_{t})\right]+\alpha^{T-1}E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[r(X_{T-1},A_{T-1})+\alpha\bar{u}^{*}(X_{T})\right]$
$\leq E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\sum_{t=0}^{T-2}\alpha^{t}r(X_{t},A_{t})\right]+\alpha^{T-1}E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\bar{u}^{*}(X_{T-1})\right]+\alpha^{T-1}\delta$
@@
$\leq E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{0},A_{0})+\alpha\bar{u}^{*}(X_{1})]-\delta+\beta\delta$
@@
$\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi(da|x_{t})\bar{\gamma}^{\delta}(dw|x_{t})\leq\bar{u}^{*}(x_{t})+\delta.$
@@
$\bar{u}^{*}(s)\leq\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}}\int_{\mathbb{W}}r(x,a)+\alpha\bar{u}^{*}(f(x,a,w))\pi^{\delta}(da|x)\psi(dw)+\delta.$ (B.6)
@@
$E_{\mu}^{\pi^{\delta},\bar{\gamma}}\left[r(X_{t},A_{t})+\alpha u^{*}(X_{t+1})\right]$
$\geq E_{\mu}^{\pi^{\delta},\bar{\gamma}}\left[\int_{\mathbb{A}}r(X_{t},a)\pi^{\delta}(da|X_{t})+\alpha\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}}\int_{\mathbb{W}}\bar{u}^{*}(f(X_{t},a,w))\psi(dw)\pi^{\delta}(da|X_{t})\right]$
$=E_{\mu}^{\pi^{\delta},\bar{\gamma}}\left[\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}}\int_{\mathbb{W}}r(X_{t},a)+\alpha\bar{u}^{*}(f(X_{t},a,w))\psi(dw)\pi^{\delta}(da|X_{t})\right]$

### v2

@@ Acknowledgments
The material in this paper is partly supported by the Air Force Office of Scientific Research under award number FA9550-20-1-0397. Support from NSF 2229012, 2312204, 2312205, 2403007, 2419564, ONR 13983111, 13983263, and 2025 New York University Center for Global Economy and Business grant is also gratefully acknowledged.
@@ Appendix A Formulations of Distributionally Robust Stochastic Control
Let $\mathbb{X},\mathbb{A},\mathbb{W}$ be Polish spaces and $(\mathbb{X},\mathcal{X}),(\mathbb{A},\mathcal{A}),(\mathbb{W},\mathcal{W})$ equip them with the Borel $\sigma$-fields. Let $\mathcal{P}(\mathcal{A})$ and $\mathcal{P}(\mathcal{W})$ be the set of probability measures on $(\mathbb{A},\mathcal{A})$ and $(\mathbb{W},\mathcal{W})$, respectively. Endow them with the topology of weak convergence; i.e. $\mu_{n}\Rightarrow\mu$ if $\int fd\mu_{n}\rightarrow\int fd\mu$ for all bounded continuous $f$. Then, $\mathcal{P}(\mathcal{A})$ and $\mathcal{P}(\mathcal{W})$ are separable, as $(\mathbb{A},\mathcal{A})$ and $(\mathbb{W},\mathcal{W})$ are separable.
@@ A.1 Admissible Policies
Notice that in this model, the history-dependent controller cannot directly use the realized action $w_{t}$ of the adversary to make its decision. This should be compared to the stochastic game settings [13] in which either player observes the action of the other and makes a decision based on such observation. However, this model can include the settings for which both players see the action of each other by considering a new state process $z_{t}=(x_{t},w_{t-1})$ and defining the state space and histories using $z_{t}$ instead of $x_{t}$.
@@ A.2 The Distributionally Robust Stochastic Control Problem
$v^{*}(\mu,\Pi,\Gamma):=\sup_{\pi\in\Pi}\inf_{\gamma\in\Gamma}v(\mu,\pi,\gamma),\quad v(\mu,\pi,\gamma):=E_{\mu}^{\pi,\gamma}\sum_{t=1}^{\infty}\alpha^{t}r(X_{t},A_{t},W_{t})$ (A.3)
@@ A.3 Dynamic Programming
In this section, we show that the solutions to the distributional robust Bellman equations (3.2) and (3.3) exists under Assumption 1, and these solutions correspond to the DRSC value (A.3).
Given the extended formulation in Appendix A, we decompose and expand the statement of Theorem 1 into the following results:
Proposition A.1.
Suppose Assumption 1 is in force. The Bellman equations (3.2) and (3.3) have unique points $u^{*}$ and $\bar{u}^{*}$, respectively within the class of uniformly continuous and non-negative functions $\mathbb{X}\rightarrow\mathbb{R}$. Moreover, $\left\|u^{*}\right\|,\left\|\bar{u}^{*}\right\|\leq\beta r_{\vee}.$
@@
Clearly, Propositon A.1, Theorem 8, and 9 implies Theorem 1. A history-dependent version of Theorem 8 is established in González-Trejo et al., [13]. In this paper, we will prove the more technically interesting Theorem 9. The proof of the rest of the Theorem 8 can be easily achieved by adapting the same proof techniques to deal with continuous state spaces in this paper to that of Theorem 2 from Wang et al., 2023b [47].
@@
Let $U_{b}(\mathbb{X})$ denote the space of uniformly bounded continuous functions on $\mathbb{X}$, which is a Banach space under the supremum norm $\left\|\cdot\right\|$.
B.1 Proof of Proposition A.1
We prove Proposition A.1 by applying the Banach fixed-point to the mapping $\mathcal{T}$ and $\overline{\mathcal{T}}$.
@@
Hence, $\left\|u^{\prime}\right\|\leq\beta r_{\vee}$.
@@ B.1.1 Proof of Lemma 2
$\leq\sup_{\phi\in\mathcal{Q}}\left|\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}r(x,a,w)+\alpha u(f(x,a,w))\phi\times\psi(da,dw)+\sup_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}-r(z,a,w)-\alpha u(f(z,a,w))\phi\times\psi(da,dw)\right|$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\sup_{\phi\in\mathcal{Q}}\sup_{\psi\in\mathcal{P}}\int_{\mathbb{A}\times\mathbb{W}}|r(x,a,w)-r(z,a,w)|+\alpha|u(f(x,a,w))-\alpha u(f(z,a,w))|\phi\times\psi(da,dw)$
$\leq\sup_{a\in\mathbb{A},w\in\mathbb{W}}|r(x,a,w)-r(z,a,w)|+\sup_{a\in\mathbb{A},w\in\mathbb{W}}\sup_{y\in\bar{B}_{f(x,a,w)}(\delta^{\prime})}\alpha|u(f(x,a,w))-u(y)|$
@@ B.2.1 Proof of Auxiliary Results for Theorem 9
Recall from A.1 that $\left\|\bar{u}^{*}\right\|\leq\beta r_{\vee}$. Define and denote the $T$-step truncated value with terminal reward $\bar{u}^{*}$ by
$v_{T}(x,\pi,\gamma):=E_{x}^{\pi,\gamma}\left[\sum_{t=0}^{T_{\epsilon}-1}\alpha^{t}r(X_{t},A_{t},W_{t})+\alpha^{T_{\epsilon}}\bar{u}^{*}(X_{T_{\epsilon}})\right].$ (B.2)
@@
$\left|v(x,\pi,\gamma)-v_{T_{\eta}}(x,\pi,\gamma)\right|$ $=\left|E_{x}^{\pi,\gamma}\left[\sum_{t=0}^{\infty}\alpha^{t}r(X_{t},A_{t},W_{t})\right]-E_{x}^{\pi,\gamma}\left[\sum_{t=0}^{T_{\eta}-1}\alpha^{t}r(X_{t},A_{t},W_{t})+\alpha^{T_{\eta}}\bar{u}^{*}(X_{T_{\eta}})\right]\right|$ (B.3)
$=\left|E_{x}^{\pi,\gamma}\left[\sum_{t=T_{\eta}}^{\infty}\alpha^{t}r(X_{t},A_{t},W_{t})-\alpha^{T_{\eta}}\bar{u}^{*}(X_{T_{\eta}})\right]\right|$
$\leq\alpha^{T_{\eta}}\left(\left|E_{x}^{\pi,\gamma}\sum_{t=T_{\eta}}^{\infty}\alpha^{t-T_{\eta}}r(X_{t},A_{t},W_{t})\right|+\left|E_{x}^{\pi,\gamma}\bar{u}^{*}(X_{T_{\eta}})\right|\right)$
@@
$\inf_{\psi\in\mathcal{P}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})$
@@
$\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta.$ (B.4)
@@
$\left\{{\psi\in\mathcal{P}:\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta}\right\}$
@@
$\left\{{h_{t}\in\mathbf{H}_{t}:\varnothing\neq\left\{{\psi\in\mathcal{P}:\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta}\right\}\cap\mathcal{D}}\right\}\in\mathcal{B}(\mathbf{H}_{t}).$
@@
$\iff$ $\exists\psi\in\mathcal{D}:\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)\leq\bar{u}^{*}(x_{t})+\delta$
$\iff$ $\inf_{\psi\in\mathcal{D}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)-\bar{u}^{*}(x_{t})-\delta\leq 0$
@@
$\lim_{k\rightarrow\infty}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w_{k}))\pi_{t}(da|h_{t})=\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})$
i.e. $w\rightarrow\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w_{k}))\pi_{t}(da|h_{t})$ is bounded continuous. Hence,
$\psi\rightarrow\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)$
@@
$c(h_{t}):=\inf_{\psi\in\mathcal{D}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\psi(dw)-\bar{u}^{*}(x_{t})-\delta$
@@
$\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi_{t}(da|h_{t})\bar{\gamma}^{\delta}_{t}(dw|h_{t})\leq\bar{u}^{*}(x_{t})+\delta.$
@@
$E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{t},A_{t},W_{t})+\alpha\bar{u}^{*}(X_{t+1})]$ $=E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{t},A_{t},W_{t})+\alpha\bar{u}^{*}(f(X_{t},A_{t},W_{t}))]$
$=E_{x}^{\pi,\bar{\gamma}^{\delta}}E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{t},A_{t}.W_{t})+\alpha\bar{u}^{*}(f(X_{t},A_{t},W_{t}))|H_{t}]$
$=E_{x}^{\pi,\bar{\gamma}^{\delta}}\int_{\mathbb{W}}\int_{\mathbb{A}}r(X_{t},a,w)+\alpha\bar{u}^{*}(f(X_{t},a,w))\pi_{t}(da|H_{t})\bar{\gamma}^{\delta}_{t}(dw|H_{t})$
@@
$v_{T}(x,\pi,\bar{\gamma}^{\delta})$ $=E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\sum_{t=0}^{T-1}\alpha^{t}r(X_{t},A_{t}.W_{t})+\alpha^{T}\bar{u}^{*}(X_{T})\right]$ (B.5)
$=E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\sum_{t=0}^{T-2}\alpha^{t}r(X_{t},A_{t},W_{t})\right]+\alpha^{T-1}E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[r(X_{T-1},A_{T-1},W_{T-1})+\alpha\bar{u}^{*}(X_{T})\right]$
$\leq E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\sum_{t=0}^{T-2}\alpha^{t}r(X_{t},A_{t},W_{t})\right]+\alpha^{T-1}E_{x}^{\pi,\bar{\gamma}^{\delta}}\left[\bar{u}^{*}(X_{T-1})\right]+\alpha^{T-1}\delta$
@@
$\leq E_{x}^{\pi,\bar{\gamma}^{\delta}}[r(X_{0},A_{0},W_{0})+\alpha\bar{u}^{*}(X_{1})]-\delta+\beta\delta$
@@
$\int_{\mathbb{W}}\int_{\mathbb{A}}r(x_{t},a,w)+\alpha\bar{u}^{*}(f(x_{t},a,w))\pi(da|x_{t})\bar{\gamma}^{\delta}(dw|x_{t})\leq\bar{u}^{*}(x_{t})+\delta.$
@@
$\bar{u}^{*}(s)\leq\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}}\int_{\mathbb{W}}r(x,a,w)+\alpha\bar{u}^{*}(f(x,a,w))\pi^{\delta}(da|x)\psi(dw)+\delta.$ (B.6)
@@
$E_{\mu}^{\pi^{\delta},\bar{\gamma}}\left[r(X_{t},A_{t},W_{t})+\alpha u^{*}(X_{t+1})\right]$
$\geq E_{\mu}^{\pi^{\delta},\bar{\gamma}}\left[\int_{\mathbb{A}}r(X_{t},a,w)\pi^{\delta}(da|X_{t})+\alpha\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}}\int_{\mathbb{W}}\bar{u}^{*}(f(X_{t},a,w))\psi(dw)\pi^{\delta}(da|X_{t})\right]$
$=E_{\mu}^{\pi^{\delta},\bar{\gamma}}\left[\inf_{\psi\in\mathcal{P}}\int_{\mathbb{A}}\int_{\mathbb{W}}r(X_{t},a,w)+\alpha\bar{u}^{*}(f(X_{t},a,w))\psi(dw)\pi^{\delta}(da|X_{t})\right]$